# Microphysical Regulation of Non-Ideal MHD in Weakly-Ionized Systems: Does the Hall Effect Matter?

### v1

@@ 2.3 The Superthermal Drift Limit
where $v_{T}$ is some characteristic slow wavespeed of the different components. Eq. 12 is equivalent to
@@ 2.3.1 Modified Collision Rates
Superthermal drift/slip speeds self-evidently violate assumption (4) in § 2.2. This means that the collision rates usually adopted, which assume negligible bulk relative drift/slip velocities of different species (so they depend on thermal velocities but not the various $u_{ij}$): $\omega_{ij}^{0}\equiv\omega_{ij}(u_{ij}\equiv|u_{i}-u_{j}|\rightarrow 0)$, are simply incorrect, even if all the other assumptions in § 2.2 held perfectly. In detail, the full collisional and inelastic couplings $\omega_{ij}$ are complicated non-linear (even non-monotonic) functions of the different relative velocities $u_{ij}$, which means they become implicit functions of ${\bf U}$, ${\bf J}$, ${\bf J}_{d}$, etc (Elmegreen, 1979; Draine & Salpeter, 1979; Pandey & Wardle, 2006; Pinto & Galli, 2008; Öberg et al., 2023). However, the leading-order behavior in the superthermal limit is straightforward. In the limit $u_{ij}\gg v_{T,i},\,v_{T,j}$, then $\omega_{ij}\propto\langle\sigma v_{ij}\rangle\propto|u_{ij}|$ increasingly moves towards the Epstein limit in which the drag force scales with the interparticle velocity. Thus, the collision rates scale with the drift/slip speed $\omega_{ij}\propto n_{j}\sigma_{ij}\Delta v_{ij}$ where $\Delta v_{ij}\propto u_{ij}=v_{{\rm drift/slip},\,ij}$. For the Epstein limit specifically as applicable to neutral hard-sphere collisions, $\omega\propto v_{T,\,ij}\,\sqrt{1+(v_{{\rm drift},\,ij}/v_{T,\,{\rm eff}})^{2}}$ with $v_{T,\,{\rm eff}}=(8/3\pi^{1/2})v_{T,\,ij}$. But this will directly modify the coefficients $\eta$, increasing $\eta_{O}$ and decreasing $\eta_{A}$ (see also Hillier 2024) proportional to the change in $\omega$.
@@
Figure 2: Ohmic, Hall, and ambipolar diffusivities $\eta_{O,H,A}$ computed for a multispecies, weakly-ionized gas with total (mostly neutral) density $n$, assuming simple fitting functions for $B$, $T$, ionization, etc. (§ 2.4.3). We compare values for slow drift+slip $v_{\rm drift},\,v_{\rm slip}\ll v_{T}$ ($\ell_{B}\gg\ell_{\rm crit}$; dotted) vs superthermal drift ($v_{\rm drift}\sim 10\,v_{T}$, $\ell_{B}\sim 0.1\,\ell_{\rm crit}^{\rm drift}$; solid) including the proposed corrections in Eq. 16 to account for enhanced scattering. We also compare $\ell_{B}\sim 0.1\,\ell_{\rm crit}$ ($v_{\rm drift,\,max}=|{\bf v}_{\rm drift}+{\bf v}_{\rm slip}|\sim 10\,v_{T}$; dashed), which has superthermal slip but not drift at low-$n$. For Hall, the lines overlap. We consider three toy models: one appropriate for protostellar core-collapse with weak $B$-fields (top); one with stronger $B$ ($B\propto n^{1/2}$ at all $n$; middle); and one with cooler temperatures and weaker ionization better matched to a circumstellar disk (bottom). This has a large effect on the absolute value of $\eta_{O,H,A}$ (“protostellar” models become hot with $T\gtrsim 1000\,{\rm K}$, at $n\gtrsim 10^{14}{\rm cm^{-3}}$, causing rapid ionization and decreasing $\eta$), but does not change the behaviors of interest here. With $\ell_{B}\gg\ell_{\rm crit}$ (sub-thermal drift), the system transitions from ambipolar to Hall to Ohmic with increasing $n$. Superthermal drift/slip greatly enhances $\eta_{O}$ and eliminates the Hall regime.
@@ 2.3.2 Anomalous Resistivity: Dropped Terms and Instabilities
Superthermal drifts also imply that many of the dropped terms associated with assumptions (2), (4)(5) can no longer be dropped. It has been well-known for decades that superthermal drifts are generically unstable on micro and mesoscales (see Kulsrud, 2005, and references therein). For systems with non-trivial chemistry (many charged species), there are a potentially infinite number of such instabilities, some relatively well-studied such as the dozen distinct current-driven instabilities (e.g. two-stream, ion-acoustic, ion-cyclotron, whistler, Buneman) reviewed in Kindel & Kennel (1971); Krall & Liewer (1971); Norman & Smith (1978); Norman & Heyvaerts (1985) or tearing (Mestel, 1968) or hybrid-mode (Davidson & Gladd, 1975) instabilities; others that have only barely been explored, such as the various novel instabilities of mixed neutral-charged systems identified in Huba (1991); Kamaya & Nishi (2000); Mamun & Shukla (2001); Nekrasov (2008b, a), or the families of intrinsically two-fluid (neutrals plus ions+electrons) drift instabilities in Tytarenko et al. (2002); Nekrasov (2009) or the charged dust+gas “resonant drag instabilities” in Squire & Hopkins (2018b); Hopkins & Squire (2018a); Seligman et al. (2019); Hopkins et al. (2020b). These instabilities grow rapidly on micro/meso-scales, producing nonlinear fluctuations that violate assumptions (2), (4)(5).
@@
Briefly, we note that while analogous instabilities exist when the neutral-charge carrier “slip” velocity (as opposed to the current drift velocity) becomes superthermal (with $v_{\rm drift}$ still subthermal; see Tytarenko et al., 2002; Squire & Hopkins, 2018b), it is less clear what their non-linear outcomes should be. This is partially for lack of study, but also because they can be stabilized on microscales by pressure effects in the neutral+ionized fluids, and some particle-based simulations of the RDI have argued that saturation can involve grains “running away” (i.e. dust-gas separation rather than stronger coupling; Moseley et al. 2019).
@@ 2.3.3 Other Assumptions
Note that even in the superthermal limit, assumption (1) from § 2.2, that displacement currents can be safely neglected, is still usually valid (producing a large displacement current would require much larger drift speeds, which we will show below is prevented by the effects discussed in § 2.3.1-2.3.2 in the regimes of interest). We briefly note that the effect of violating (1) would be to decouple $B$ so it detaches and propagates at $c$, akin to an even much larger anomalous resistivity. In (2), the neglect of the battery terms, which requires $a_{\rm ext}\ll\omega_{i}v_{\rm drift}$, is not invalidated by superthermal drift (so long as it was already valid with sub-thermal drift, its validity should only be strengthened by superthermal drift). Assumption (3), while common, is completely unnecessary for Eq. 8 – one can simply include the appropriate coefficients $\omega^{{\rm ion/rec},0}$ and $\omega^{\rm coll}_{-+}$ in $\omega_{i}^{0}$.
@@ 2.4.1 Corrected Collision Rates
Nonetheless, the discussion above suggests that we can capture the leading-order effects from § 2.3 by (1) multiplying all collision rates by $\sim[1+(v_{\rm drift,\,slip}/v_{T})^{2}]^{1/2}$, so they correctly capture the leading-order superthermal corrections to the drag law, and (2) adding an approximate anomalous resistivity as an additional effective collision rate, which “turns on” at $v_{\rm drift}>v_{T}$ and therefore self-regulates the system to $v_{\rm drift}\lesssim v_{T}$. Noting the dependence of the various $\eta$ on the collision frequencies in § 2.1, this becomes:
$\eta_{O}^{\rm eff}$ $\rightarrow\left(\eta_{O}^{0}+\eta_{O}^{\rm an}\,\Theta\left[\frac{v_{\rm drift}}{v_{T}}\right]\right)\,\left[1+\left(\frac{v_{\rm drift}}{v_{T}}\right)^{2}\right]^{1/2}\ ,$
$\eta_{O}^{\rm an}$ $\equiv\frac{c^{2}m_{-}}{4\pi|q_{-}|^{2}\bar{n}_{-}}\omega^{\rm an}\ ,$
@@
$\eta_{A}^{\rm eff}$ $\rightarrow\eta_{A}^{0}\,\left[1+\left(\frac{v_{\rm drift,\,max}}{v_{T}}\right)^{2}\right]^{-1/4}\ ,$ (16)
@@
where $\Theta(x)$ can be the Heaviside step function or any similar function (e.g. $e^{-1/x}$), and $v_{\rm drift,\,max}$ is defined from Eq. 9 using $\eta_{A}^{0}$ (i.e. from the usual coefficient estimated without including velocity-dependence). Note (as justified more quantitatively in § A), the dependence of $\eta_{O}$ on $v_{\rm drift}$ (as the relevant terms in the collision rates depend more strongly on $v_{\rm drift}$ in the limit where this is large), and the $1/4$ slope in $\eta_{A}^{\rm eff}$ arises because of the implicit non-linear dependence of $v_{\rm slip}$ on $\eta_{A}$ (hence the collision rate $\omega_{\rm coll}^{\rm eff,\,A}$ in Eq. 9; see § A and Hillier 2024). It is not necessary to drop inelastic reactions in the $\eta^{0}$ terms if they are important. Motivated by the extensive studies discussed § 2.3, we take $\omega^{\rm an}$ to be the faster of either the gyro or plasma frequencies, e.g. $\omega^{\rm an}\sim(\bar{\Omega}_{-}^{2}+\omega_{p-}^{2})^{1/2}$, but we will show that the exact value does not matter so long as it reflects a sufficiently fast frequency of the system.
@@
Prescriptions similar to this are already widely used in laboratory, Solar, and ionospheric plasma physics applications (see references in § 2.3.2 and Somov & Oreshina 2000; Roussev et al. 2002; Ni et al. 2007; Færder et al. 2023, for examples). Likewise, it is directly analogous to widely adopted treatments of similar problems in astrophysical applications of thermal conduction and viscosity, where naively applying the “classical” conduction/viscosity equations (Spitzer & Härm, 1953; Braginskii, 1965) can produce sharp gradients that imply unphysically superthermal drift velocities (of e.g. thermal electrons or ions in the “saturated” limit; e.g. Spitzer 1962; Chapman & Cowling 1970; Cowie & McKee 1977). It has also been widely recognized that for high-$\beta^{\rm plasma}$ plasmas, the maximum anisotropy of the velocity distribution function (which can be considered similarly to a maximum drift speed) is limited by microscopic instabilities like the whistler, mirror, and firehose, which are activated above certain thresholds (Schekochihin2008b; Komarov et al. 2014, 2016; Komarov et al. 2018; Kunz et al. 2014; Squire et al. 2016; Squire et al. 2017a; Squire et al. 2017b). These are similarly approximated on macroscopic scales by multiplying the “classical” coefficients by corrections designed to represent some anomalous scattering that limits the velocity anisotropy (see e.g. Riquelme et al., 2016; Roberg-Clark et al., 2016, 2018; Su et al., 2017; Squire et al., 2017c; Hopkins et al., 2020a). We are simply pointing out that similar terms should apply in the other astrophysical cases of interest here (as was also argued by authors like Norman & Heyvaerts 1985), generalizing them to cases with e.g. appreciable neutrals and non-electron-ion plasmas, while also accounting for similar the relevant terms in the ambipolar drift.

### v2

@@ 2.3 The Superthermal Drift Limit
where $v_{T}$ is some characteristic slow wavespeed of the different components. Specifically, this is explicitly part of assumption (4) (§ 2.3.1; see also Appendix A), and implicitly (for reasons discussed in more detail below) part of assumptions (2) & (5) (assumptions (1) and (3), on the other hard, are not so directly affected by violations of Eq. 12, though we discuss what might happen if they break down below as well). From Eq. 9 for $v_{\rm drift}$ and $v_{\rm slip}^{\pm}$, we see that Eq. 12 is equivalent to:
@@ 2.3.1 Modified Collision Rates
Superthermal drift/slip speeds self-evidently violate assumption (4) in § 2.2. This means that the collision rates usually adopted, which assume negligible bulk relative drift/slip velocities of different species (so they depend on thermal velocities but not the various $\Delta u_{ij}\equiv|{\bf u}_{i}-{\bf u}_{j}|$): $\omega_{ij}^{0}\equiv\omega_{ij}(\Delta u_{ij}\rightarrow 0)$, are simply incorrect, even if all the other assumptions in § 2.2 held perfectly.
@@
If we retain all other assumptions from § 2.2, but allow the various $\omega_{ij}$ to be functions of $\Delta u_{ij}$ as above, then Eqs. 8 & 9 apply but with modified coefficients $\eta\neq\eta^{0}$. In detail, the full collisional and inelastic couplings $\omega_{ij}$ are complicated non-linear (even non-monotonic) functions of the different relative velocities $\Delta u_{ij}$, which means they become implicit functions of $\bf U$}, $\bf J$}, ${\bf J}_{d}$, (Elmegreen, 1979; Draine & Salpeter, 1979; Pandey & Wardle, 2006; Pinto & Galli, 2008; Öberg et al., 2023). This means the coefficients $\eta$ must be calculated numerically, which we show for simple three and five-species models in § A. However, in § B.1 we use our three-species derivation (§ 2.1) to show that the leading-order behavior in the superthermal limit can be approximated with simple scalings. Briefly, in the limit $\Delta u_{ij}\gg v_{T,i},\,v_{T,j}$, then $\omega_{ij}\propto\langle\sigma v_{ij}\rangle\propto\Delta u_{ij}$ increasingly moves towards the Epstein limit in which the drag force scales with the interparticle velocity. Thus, the collision rates scale with the drift/slip speed $\omega_{ij}\sim n_{j}\langle\sigma v_{ij}\rangle^{0}\sqrt{1+\Delta u^{2}_{ij}/v_{T,\,ij}^{2}}$ where $\Delta u_{ij}\sim v_{{\rm drift/slip},\,ij}$. If we simply multiply the collision rates by some factor $f_{\omega}$, we see from Eq. 5 that $\eta_{O}\propto\omega_{O}$ scales $\propto f_{\omega}$, while $\eta_{H}$ is independent of $f_{\omega}$ and $\eta_{A}\propto\omega_{A}^{-1}\propto f_{\omega}^{-1}$. Thus after consideration of which species and relative speeds $\Delta u_{ij}$ appear in each, we obtain the approximate Eq. 25: $\eta_{O}\sim\eta_{O}^{0}\,[1+(|v_{\rm slip,\,eff}|^{2}+|{\bf v}_{\rm drift}^{0}|^{2})/v_{T,\,\rm eff}^{2}]^{1/2}$, $\eta_{H}\sim\eta_{H}^{0}$, $\eta_{A}\sim\eta_{A}^{0}\,[1+|v_{\rm slip,\,eff}|^{2}/v_{T,\,\rm eff}^{2}]^{1/2}$, where $v_{\rm slip,\,eff}$ is defined there.
@@
Figure 2: Ohmic, Hall, and ambipolar diffusivities $\eta_{O,H,A}$ computed for a multispecies, weakly-ionized gas with total (mostly neutral) density $n$. We compare values for slow drift+slip $v_{\rm drift},\,v_{\rm slip}\ll v_{T}$ ($\ell_{B}\gg\ell_{\rm crit}$; dotted) vs superthermal drift ($v_{\rm drift}\sim 10\,v_{T}$, $\ell_{B}\sim 0.1\,\ell_{\rm crit}^{\rm drift}$; solid) including the proposed corrections in Eq. 16 to account for enhanced scattering. We also compare $\ell_{B}\sim 0.1\,\ell_{\rm crit}$ ($v_{\rm drift,\,max}=|{\bf v}_{\rm drift}+{\bf v}_{\rm slip}|\sim 10\,v_{T}$; dashed), which has superthermal slip but not drift at low-$n$. For Hall, the lines overlap. We consider three models for $B$, $T$, ionization, etc. from NICIL (§ 2.4.3): a protostellar collapse model with non-ideal MHD (top), a barytropic equation-of-state model with stronger $B$ (middle), and a circumstellar disk midplane model with lower $T$ and weaker ionization (bottom). The models change the value of $\eta_{O,H,A}$ (e.g. collapse models become hot at $n\gtrsim 10^{14}{\rm cm^{-3}}$, causing rapid ionization and decreasing $\eta$), but do not change the systematic offsets of interest here. With $\ell_{B}\gg\ell_{\rm crit}$ (sub-thermal drift), the system transitions from ambipolar to Hall to Ohmic with increasing $n$. Superthermal drift/slip greatly enhances $\eta_{O}$ and eliminates the Hall regime.
@@ 2.3.2 Anomalous Resistivity: Dropped Terms and Instabilities
Superthermal drifts also imply that many of the dropped terms associated with assumptions (2), (4) & (5) in § 2.2 can no longer be dropped. It has been well-known for decades that superthermal drifts are generically unstable on micro and mesoscales (see Kulsrud, 2005, and references therein). For systems with non-trivial chemistry (many charged species), there are a potentially infinite number of such instabilities, some relatively well-studied such as the dozen distinct current-driven instabilities (e.g. two-stream, ion-acoustic, ion-cyclotron, whistler, Buneman) reviewed in Kindel & Kennel (1971); Krall & Liewer (1971); Norman & Smith (1978); Norman & Heyvaerts (1985) or tearing (Mestel, 1968) or hybrid-mode (Davidson & Gladd, 1975) instabilities; others that have only barely been explored, such as the various novel instabilities of mixed neutral-charged systems identified in Huba (1991); Kamaya & Nishi (2000); Mamun & Shukla (2001); Nekrasov (2008b, a), or the families of intrinsically two-fluid (neutrals plus ions+electrons) drift instabilities in Tytarenko et al. (2002); Nekrasov (2009) or the charged dust+gas “resonant drag instabilities” in Squire & Hopkins (2018b); Hopkins & Squire (2018a); Seligman et al. (2019); Hopkins et al. (2020b). These instabilities grow rapidly on micro/meso-scales, producing nonlinear fluctuations that violate assumptions (2), (4) & (5).
@@
While the details of the anomalous scalings are uncertain and complicated, as shown in § B.2, if we insert any of the heuristic scalings above into our three-species model in § 2.1 (e.g. modify the anisotropic pressure terms, account for various effective Reynolds-averaged terms from fluctuations, or allow for effective anomalous collisional damping), we obtain the same qualitative scaling. In each case Eq. 8 remains valid with the ambipolar and Hall terms unmodified to leading order ($\eta_{H}\rightarrow\eta_{H}^{0}$, $\eta_{A}\rightarrow\eta_{A}^{0}$), but $\eta_{O}\rightarrow\eta_{O}^{0}+\eta_{O}^{\rm an}$ (i.e. the anomalous term adds linearly in $\eta_{O}$), where dimensionally $\eta_{O}\sim(c^{2}\,m_{-}/4\pi\,q_{-}^{2}n_{-})\,\omega^{\rm an}\,(1+v^{2}_{\rm drift}/v^{2}_{T})^{2}$ for $v_{\rm drift}\gtrsim v_{T}$.
Briefly, we note here and in § B.2 that while analogous instabilities exist when the neutral-charge carrier “slip” velocity (as opposed to the current drift velocity) becomes superthermal (with $v_{\rm drift}$ still subthermal; see Tytarenko et al., 2002; Squire & Hopkins, 2018b), it is less clear what their non-linear outcomes should be. This is partially for lack of study, but also because they can be stabilized on microscales by pressure effects in the neutral+ionized fluids, and some particle-based simulations of the RDI have argued that saturation can involve grains “running away” (i.e. dust-gas separation rather than stronger coupling; Moseley et al. 2019).
@@ 2.3.3 Other Assumptions
Note that even in the superthermal limit, assumption (1) from § 2.2, that displacement currents can be safely neglected, is still usually valid (producing a large displacement current would require much larger drift speeds, which we will show below is prevented by the effects discussed in § 2.3.1-2.3.2 in the regimes of interest). We briefly note that the effect of violating (1) would be to decouple $B$ so it detaches and propagates at $c$, akin to an even much larger anomalous resistivity. In (2), the neglect of the battery terms, which requires $a_{\rm ext}\ll\omega_{i}v_{\rm drift}$, is not invalidated by superthermal drift (so long as it was already valid with sub-thermal drift, its validity should only be strengthened by superthermal drift). Assumption (3), while common, is not strictly necessary for Eq. 8 – one can simply include the appropriate coefficients $\omega^{{\rm ion/rec},0}$ and $\omega^{\rm coll}_{-+}$ in $\omega_{i}^{0}$.
@@ 2.4.1 Corrected Collision Rates
Our discussion in § 2.3 & B argues that we can capture the leading-order effects of superthermal drift in our derivation (§ 2.1) by (1) modifying the effective collision rates $\omega_{\rm coll}^{0}$ by multiplying by the superthermal drift/slip dependent corrections (§ 2.3.1, B.1), and (2) adding an appropriate anomalous resistivity which “turns on” at $v_{\rm drift}>v_{T}$ (§ 2.3.2, B.2). So our effective coefficients become:
$\eta_{O}^{\rm eff}$ $\rightarrow\eta_{O}^{0}\,\left[1+\frac{v_{\rm slip,\,eff}^{2}+|{\bf v}_{\rm drift}^{0}|^{2}}{v_{T}^{2}}\right]^{1/2}+\eta_{O}^{\rm an}\,\Theta\left[\frac{v_{\rm drift}}{v_{T}}\right]\ ,$
$\eta_{O}^{\rm an}$ $\equiv\frac{c^{2}m_{-}}{4\pi|q_{-}|^{2}\bar{n}_{-}}\omega^{\rm an}\left[1+\frac{|{\bf v}_{\rm drift}^{0}|^{2}}{v_{T}^{2}}\right]^{1/2}\ ,$
@@
$\eta_{A}^{\rm eff}$ $\rightarrow\eta_{A}^{0}\,\left[1+\frac{v_{\rm slip,\,eff}^{2}}{v_{T}^{2}}\right]^{-1/2}\ ,$ (16)
@@
where $v_{\rm slip,\,eff}^{2}/v_{T}^{2}\equiv-1/2+\sqrt{1/4+|{\bf v}_{\rm slip,\,\bot}^{0}|^{2}/v_{T}^{2}}$ and $v_{\rm drift,\,max}^{2}\equiv v_{\rm slip,\,eff}^{2}+|{\bf v}_{\rm drift}^{0}|^{2}$ (with ${\bf v}_{\rm slip,\,\bot}^{0}$ and ${\bf v}_{\rm drift}^{0}$ defined in Eq. 9, using $\eta_{A}^{0}$ – i.e. from the usual coefficients without velocity-dependence). Here $\Theta(x)$ can be the Heaviside step function or any similar function (e.g. $e^{-1/x}$). It is not necessary to drop inelastic reactions in the $\eta^{0}$ terms if they are important. Motivated by the extensive studies discussed § 2.3, we take $\omega^{\rm an}$ to be the faster of either the gyro or plasma frequencies, e.g. $\omega^{\rm an}\sim(\bar{\Omega}_{-}^{2}+\omega_{p-}^{2})^{1/2}$, but we will show that the exact value does not matter so long as it reflects a sufficiently fast frequency of the system. Eq. 16 is approximate: in § A we discuss how more accurate correction terms can be obtained numerically, but show Eq. 16 captures their important asymptotic scalings.
Prescriptions akin to this are already used in laboratory, Solar, and ionospheric plasma physics applications (see references in § 2.3.2 and Somov & Oreshina 2000; Roussev et al. 2002; Ni et al. 2007; Færder et al. 2023, for examples). However, Eq. 16 expands upon these (based on our discussion above) by (1) including Ohmic, Hall, and ambipolar terms (as opposed to just Ohmic); (2) including the drift-dependent corrections to the $\eta_{0}$ terms (usually ignored); (3) generalizing $\eta^{\rm an}$ to allow both weakly ($\Omega_{e}<\omega_{pe}$, assumed in those papers) and strongly ($\Omega_{e}>\omega_{pe}$) magnetized limits; (4) allowing for mostly-neutral gas (as compared to just fully-ionized fluids); (5) generalizing (§ 2.4.2 & A below) to multi-species cases; and (6) providing criteria for the critical drift/slip speeds in different media.
@@
Likewise, Eq. 16 is analogous to widely adopted treatments of similar problems in astrophysical applications of thermal conduction and viscosity, where naively applying the “classical” conduction/viscosity equations (Spitzer & Härm, 1953; Braginskii, 1965) can produce sharp gradients that imply unphysically superthermal drift velocities (of e.g. thermal electrons or ions in the “saturated” limit; e.g. Spitzer 1962; Chapman & Cowling 1970; Cowie & McKee 1977). It has also been widely recognized that for high-$\beta^{\rm plasma}$ plasmas, the maximum anisotropy of the velocity distribution function (which can be considered similarly to a maximum drift speed) is limited by microscopic instabilities like the whistler, mirror, and firehose, which are activated above certain thresholds (Schekochihin et al. 2008; Komarov et al. 2014, 2016; Komarov et al. 2018; Kunz et al. 2014; Squire et al. 2016; Squire et al. 2017a; Squire et al. 2017b). These are similarly approximated on macroscopic scales by multiplying the “classical” coefficients by corrections designed to represent some anomalous scattering that limits the velocity anisotropy (see e.g. Riquelme et al., 2016; Roberg-Clark et al., 2016, 2018; Su et al., 2017; Squire et al., 2017c; Hopkins et al., 2020a). We are simply pointing out that similar terms should apply in the other astrophysical cases of interest here (as was also argued by authors like Norman & Heyvaerts 1985), generalizing them to cases with e.g. appreciable neutrals and non-electron-ion plasmas, while also accounting for similar the relevant terms in the ambipolar drift.